OpenStudy (anonymous):

Use a triple integral to find the volume of the given solid: The solid bounded by the cylinder y = x^2 and the planes z = 0 , z = 4 and y =9 Please explain step by step.

OpenStudy (dumbcow):

this problem can actually be done using only 1 integral but i will show you the longer triple integral first determine limits on variables: z -> 0 to 4 y -> 0 to 9 x -> -sqrty to sqrty y = x^2 --> x = +-sqrt(y) \[V = \int\limits_{0}^{4}\int\limits_{0}^{9}\int\limits_{-\sqrt{y}}^{\sqrt{y}} dx dy dz\]

OpenStudy (anonymous):

would you it be: \[V \int\limits_{0}^{4} \int\limits_{0}^{9} \int\limits_{-\sqrt{9}}^{\sqrt{9}} x^2 dx dydz \]

OpenStudy (anonymous):

integrating to get x^3/3 evaluated at (-sqrt 9 to sqrt 9)

OpenStudy (dumbcow):

i don't think so...i guess i could be wrong setting up the triple integral but i know the answer is 144 area inside parabola y=x^2 from -3 to 3 is 36 36*4 = 144

OpenStudy (anonymous):

It is 144 how would i evaluate the integral if not that way?

OpenStudy (dumbcow):

\[4\int\limits_{-3}^{3}(9-x^{2}) dx\]

OpenStudy (anonymous):

where does the 4 come from? and that would give u = 9-x^2 du = -x^(3)/3dx ad im unsure where to go from there

OpenStudy (dumbcow):

height of cylinder is 4, z goes from 0 to 4 no substitution needed, just use power rule (its a polynomial)

OpenStudy (anonymous):

oh its just 9x- x^(3)/3? evaluated from 2 to -2 and then moving onto dydz with 4 in the front?